mr wasike wants to invest a certain portion of his lump sum retirement benefits in order to accumulate 300000 shillings in five years.How much should he invest if the money will earn 13% per annum?Homw much interest will be earned over 5years
Answer by stanbon(75887)

mr wasike wants to invest a certain portion of his lump sum retirement benefits in order to accumulate 300000 shillings in five years.How much should he invest if the money will earn 13% per annum?Homw much interest will be earned over 5years
------
Assuming simple interest,
A(t) = P(1+rt)
300000 = P(1+0.13*5)
P = 300000/(1.65)
P = 181818.81 (amt to invest)
---
300,000-181818.81 = 118181.82 (amt of interest)
----------------
Cheers,
Stan H.